Tea worth ₱25.50 per pound is to be mixed
with tea worth ₱29.40 per pound. How many pounds of each should be used to obtain 25 pounds of a blend worth ₱27.00 per pound?

Answer by richwmiller(17219)

25*27=675
x+y=25,
25.5*x+29.4*y=675
x=25-y
25.5*(25-y)+29.4*y=675
637.5-25.5y+29.4*y=675
3.9*y=37.5
y=9.61538462
x=25-y
x=15.3846154 at 25.50 and y=9.61538462 at 29.40
check
15.3846154+9.61538462=25
25.5*x+29.4*y=675
25.5*15.3846154+29.4*9.61538462=675
392.307692+282.692308=675
675.0=675
ok